You’ve collected the following information from your favorite financial website.
52-Week Price Stock (Div) Div
Yld %
PE
Ratio
Close
Price
Net
Chg
Hi Lo
64.60        47.80          Abbott 1.12 1.9       235.6     62.91     −.05    
145.94        70.28          Ralph Lauren 2.50 1.8       70.9     139.71    .62    
171.13        139.13          IBM 6.30 4.3       23.8     145.39     .19    
91.80        71.96          Duke Energy 3.56 4.9       17.6     74.30     .84    
113.19        96.20          Disney 1.68 1.7       15.5     ??       .10    
Find the quote for the Duke Energy. Assume that the dividend is constant.
a. What was the highest dividend yield over the past year? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
b. What was the lowest dividend yield over the past year? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

1.
Highest dividend yield occurs for lowest price=3.56/71.96=4.95%

2.
Lowest dividend yield occurs for highest price=3.56/91.80=3.88%